### v6

# Brain-inspired Large Language Model

Abstract
We introduce BriLLM, a brain-inspired large language model that redefines the foundations of generative language modeling. Departing from Transformer architectures, GPT frameworks, and traditional input-output constrained paradigms, BriLLM is built on the Signal Fully-connected flowing (SiFu) mechanism—a directed graph-based neural network design that enables full interpretability across all nodes, in contrast to conventional models limited to input-output interpretability. In this framework, tokens are represented as graph nodes, with signal flows—either randomly initialized or user-defined—propagating along paths following a "least resistance" principle. The next token to be generated emerges as the target of this signal flow. Theoretically, BriLLM supports infinitely long $n$-gram modeling, with model size decoupled from input and prediction length. Its signal propagation dynamics mimic human-like cognitive patterns, enabling recall activation and inherent multi-modal compatibility. We release initial Chinese and English BriLLM versions (4000 tokens, 32-dimensional nodes, 32-token sequence prediction capacity) with sizes  2B and  1B parameters, respectively, achieving performance comparable to GPT-1.  [^1] Code at: https://github.com/brillm05/BriLLM0.5  [^2] Models at: https://huggingface.co/BriLLM/BriLLM0.5
1 Introduction
Large language models (LLMs) have ignited unprecedented progress toward artificial general intelligence (AGI). Yet, state-of-the-art LLMs—rooted in Transformer architectures and GPT (Generative Pre-trained Transformer) training schemes—face fundamental limitations that may preclude them from achieving true AGI. Their scalability is constrained by computational bottlenecks, and their interpretability remains superficial, while their operational principles bear little resemblance to the human brain: a naturally evolved intelligence system refined over millions of years that offers a blueprint for AGI.
The Transformer (Vaswani et al., 2017) has become the backbone of modern LLMs, delivering remarkable generalization across tasks and scaling to increasingly sophisticated capabilities. However, the attention mechanism’s quadratic computational complexity with respect to sequence length creates intractable inefficiencies for long inputs. While efforts to develop linear attention variants or RNN-like Transformers aim to balance performance and cost, they merely mitigate rather than resolve this core bottleneck, as they retain attention-based architectures.
Compounding these challenges, Transformers suffer from limited parameter-level interpretability due to opaque self-attention dynamics and complex parameter interactions, functioning as "black boxes." Moreover, when deploying LLMs, GPT requires a quadratic increase in model size as input context length grows linearly. What is more, multi-modal large models based on Transformers or GPT depend heavily on multi-modal aligned data, making it far harder to build embodied intelligent systems that resemble human-like ones.
To address these critical limitations—which are not unique to Transformers or GPT but inherent to traditional machine learning—we rethink machine learning from its fundamental principles, integrating brain-inspired computational ideas. This is a straightforward line of reasoning: the brain, the seat of human intelligence, must provide key insights for overcoming the limitations of current machine learning frameworks as we move toward AGI. This rethinking has led to a novel machine learning mechanism termed Signal Fully-connected Flowing (SiFu) in this paper. SiFu learning is defined over a fully connected directed graph where all nodes are interpretable. Signaling in SiFu replaces the core prediction operations in traditional machine learning.
Building on the SiFu mechanism, we present BriLLM: a language model inspired by neural signal propagation in the brain. Unlike attention-based architectures, BriLLM functions on a bidirectional graph with nodes and edges, where each node (a neuron layer) represents a token, and edges—modeled as fully connected networks—encode token relationships. Signal propagation adheres to energy dynamics: tensors flow through the graph, with the next token determined by the path that maximizes signal energy.
Table 1 compares the two frameworks of existing machine learning and SiFu learning.
Table 1: Machine learning vs SiFu learning frameworks
GPT-LLM BriLLM
$\uparrow$ various "large" models various brain-like "large" models
GPT-LLM BriLLM
Transformer, GPT
Deep Learning SiFu Learning
Machine Learning
2 SiFu Mechanism
The human brain’s representation and processing of information—particularly language—have been characterized in Huth et al. (2016): semantic features of linguistic elements are distributed across the cerebral cortex, with distinct cortical regions encoding specific semantic content, and this organization exhibits inter-individual consistency. This established neural basis of language contrasts sharply with the representational and processing mechanisms of current GPT-LLMs, which only maintain interpretable semantic mappings for linguistic elements at their input and output interfaces. In contrast, the human brain encodes semantics directly within distributed cortical regions, rendering it a fully interpretable "large model" if conceptualized as a language-processing system.
Furthermore, the brain differs fundamentally from GPT-LLMs and all existing machine learning models in how it regulates input-output dynamics. The latter process a discrete input stream through opaque internal computations to generate a predictive output stream, with input and output restricted to dedicated interfaces. The brain, however, almost certainly operates via a distinct mechanism. While experimental and ethical constraints limit our understanding of its detailed operational processes (i.e., thinking), the distributed cortical encoding of semantics implies that language-based cognition requires all semantic units (e.g., vocabulary) to function as both inputs and outputs. Consequently, the cortical regions subserving these semantic representations must act as bidirectional nodes for information flow, serving as both input and output sites. Conceptualizing the brain as a machine learning model thus necessitates a mechanism enabling all components to operate as simultaneous input and output interfaces. This requirement leads us to hypothesize a dynamic signaling system—likely electrophysiological signals such as EEG—that can activate these distributed semantic regions (or model components) at any time.
Inspired by the supposed neural signaling in the brain, we introduce Signal Fully-connected Flowing (SiFu)—a novel input-output control mechanism over directed graphs that forms the core of BriLLM. This mechanism comprises two core components: (1) a fully connected directed graph with bidirectional node connections; (2) a dynamic weighting system that modulates signal strength based on node correlations. As shown in Figure 2, the SiFu model is a graph with sparsely activated nodes that transmit nominal signals through tensors. Each node (ideally a neuronal layer) represents a concept or word (e.g., nouns, verbs), while edges model inter-node relationships.
Figure 1: The schematic illustration of SiFu mechanism.
We first formalize the generative machine learning task addressed here. For language models, generative prediction—also termed autoregressive prediction—originates from the classic $n$-gram language modeling framework. Such tasks aim to predict the next token $w_{i}$ from the preceding sequence $w_{1},w_{2},\ldots,w_{i-1}$. In deep learning implementations, this requires training a model $M$ such that
$\text{Id}(w_{i})=M_{\theta}(e(w_{1}),\ldots,e(w_{i-1})),$
where $\text{Id}()$ denotes the token’s output representation (typically a one-hot vector), $e()$ denotes the input representation (commonly called word vectors), and $\theta$ is the set of parameters to be learned by $M$. Notably, such models must integrate prior outputs into the context for prediction, giving rise to ’attention’ mechanisms—common to both RNN and Transformer architectures. Accordingly, the attention-augmented form becomes
$\text{Id}(w_{j})=M_{\theta}(e(w_{1}),\ldots,e(w_{i-1});\text{Attention}(w_{i},w_{i-1},\ldots,w_{j-1})).$
In these models, only the input sequence $w_{1},\ldots,w_{i-1}$ and output $w_{i}$ are directly interpretable; the model $M$ and its parameters $\theta$ require dedicated analysis to elucidate their roles in the learning process. Moreover, model size scales with input context length, as $M$ must process the entire sequence $w_{1},\ldots,w_{i-1}$.
To address these limitations, the SiFu learning mechanism introduces a redesigned model and prediction paradigm. SiFu is defined as a directed fully connected graph $G=\{V,E\}$, where $V$ (nodes) and $E$ (edges) form a fully connected structure by default. Each node in $V$ is explicitly mapped to a token in the vocabulary: for a vocabulary $\{w_{1},w_{2},\ldots,w_{n}\}$ (with size $n$), the node set is $V=\{v_{1},v_{2},\ldots,v_{n}\}$. This design inherently renders the model interpretable, as each node’s meaning is user-defined—though node semantics (i.e., associated tokens) are distinct from the graph’s structural properties. We assume nodes and edges have specific, learnable topological structures and parameters $\theta_{V}$ (node’s hyperparameters and parameters) and $\theta_{E}$ (edge’s hyperparameters and parameters) to govern tensor flow $r$, which propagates across edges and through nodes.
Signal transformation at nodes and edges follows: $r^{\prime}=r\oplus v_{i,\theta_{V}}$ after passing through node $v_{i}$, and $r"=r\otimes e_{ij,\theta_{E}}$ after traversing edge $e_{ij}$, where $\oplus$ and $\otimes$ denote predefined tensor operations for node and edge interactions, respectively. For simplicity, autoregressive prediction focuses on correctly forecasting the next token. In SiFu, given an input sequence $w_{1},\ldots,w_{i-1}$ (mapped to nodes $v_{1},\ldots,v_{i-1}$), a signal tensor $r$ initiates at $v_{1}$, propagates sequentially through these input nodes to $v_{i-1}$, and then continues onward. The next predicted node is determined by maximizing signal energy: the correct node $v_{i}$ is the one where $r$ attains the highest energy $\|r\|$ (with $\|\cdot\|$ denoting a custom tensor energy function), such that
$v_{i}=\arg\max_{v^{\prime}}\left\|r\oplus v_{1}\otimes e_{12}\oplus v_{2}\ldots\oplus v^{\prime}\right\|.$
However, to predict the token $v_{i}$, relying solely on the information from the signal tensor passed from the last preceding node $v_{i-1}$, while ignoring all previous nodes, can lead to significant prediction bias. Therefore, like traditional generative models, we also use a mechanism of ’attention’ to combine the signal tensors output by each node before the node to be predicted. Based on the energy score of the fused signal tensor, we make the final prediction, and thus the predictive formula that can ’attend’ to the information of all nodes in the previous input sequence is
$v_{i}=\arg\max_{v^{\prime}}\sum_{k=1}^{i-1}\left\|\alpha_{k}*(r\oplus v_{1}\otimes e_{12}\oplus v_{2}\ldots\otimes e_{k-1,k}\oplus v_{k}\otimes e_{k,v^{\prime}}\oplus v^{\prime})\right\|.$
where $\alpha_{k}$ represents a learnable weight for signaling from $v_{k}$ to $v^{\prime}$.
This defines SiFu’s prediction process (also called inference or decoding). During training, model parameters may simply optimized to reward correct node predictions and penalize incorrect ones.
In summary, signal propagation is governed by energy dynamics: signals strengthen (maximizing energy) along ’correct’ paths, and the next node is activated only if it maximizes signal energy, fostering competition between potential paths. In details, given input tokens, a signal initiates at the starting node and propagates sequentially. Outputs correspond to nodes along the path with maximal signal energy relative to alternatives. For example, Figure 2 illustrates "dog $\to$ love $\to$ meat" as the highest-energy path. During training (Figure 2), the ground-truth sequence must yield the highest energy. For "love $\to$ meat," for instance, energy tensors are computed for the target ("meat") and negative samples ("apple," "iron"), with the loss function maximizing energy for "meat" while minimizing it for negatives.
(a) Forward
(b) Optimization
Figure 2: An illustration of SiFu Directed Graph (Numbers by the node denote energy scores).
Overall, SiFu’s design as a directed fully connected graph with signal propagation confers two key advantages:
  1. 1.
    Inherent full interpretability: User-defined entities (concepts, tokens, or interpretable units) map directly to specific graph nodes;
  2. 2.
    Unbounded contextual capacity: Prediction is framed as signal propagation through node activations. Because signals propagate freely across nodes, sequence prediction naturally supports arbitrarily long contexts without increasing model size.
3 BriLLM Formulation
BriLLM implements the SiFu mechanism for language modeling, as shown in Figure 4. Each vocabulary token corresponds to a node, defined by a GeLU-activated neuron layer with a bias $b\in\mathbb{R}^{d_{node}}$, where $d_{node}$ is the node dimension (number of neurons per node). Edges between nodes $u$ and $v$ are modeled as fully connected matrices $W_{u,v}\in\mathbb{R}^{d_{node}\times d_{node}}$, with $W_{u,v}$ and $W_{v,u}$ enabling bidirectional signaling.
Signal propagation begins with an initial tensor:
$e_{0}=[1,1,\dots,1]^{\top}\in\mathbb{R}^{d_{node}}$ (1)
For a training sequence $u_{1},...,u_{L-1},v_{predict}$, signal $e_{i+1}\in\mathbb{R}^{d_{node}}$ when propagating from $u_{i}$ to $u_{i+1}$ is computed as:
$e_{{i+1}}=\left\{\begin{aligned} {\rm GeLU}(W_{u_{i},u_{i+1}}e_{i}+b_{u_{i},u_{i+1}}+PE_{i})\;\;{\rm if}\;\;i>0&\\ {\rm GeLU}(e_{0}+b_{u_{1}}+PE_{0})\;\;{\rm if}\;\;i=0\end{aligned}\right$.
where PE denotes sine-cosine positional encoding. Edge-specific biases $b_{u_{i},u_{i+1}}$ modulate signaling, with $b_{u_{1}}$ (edge-independent) initiating sequences.
To predict the next token, an aggregated signal tensor $\mathcal{E}_{i}\in\mathbb{R}^{d_{node}}$ is computed via a weighted sum of previous signals using learnable weights $\alpha\in\mathbb{R}^{L-1}$:
$\mathcal{A}$ $=\mathop{{\rm softmax}}(\alpha_{1:L-1})$ (2)
$\mathcal{E}_{L-1}$ $=\sum_{k=1}^{L-1}\mathcal{A}_{k}e_{k},$ (3)
where $L$ is sequence length, and $\mathcal{A}$ represents softmax-normalized weights. This allows the model to "attend" to all preceding tokens, not just the immediate predecessor. Note here $\mathcal{A}$ represents a rank-1 tensor weight for attending each previous signal. However, we empirically discovered higher ranked weighting tensor may further strengthen the long sequence prediction of BriLLM.
The final signal tensor for predicting token $v$ is:
$E_{u,v}={\rm GeLU}(W_{u_{L-1},v}\mathcal{E}_{L-1}+b_{u_{L-1},v}+PE_{L-1}),$
During inference, the predicted token $v_{predict}$ maximizes the L2 norm of this signal tensor:
$v_{predict}$ $=\mathop{\arg\max}_{v}\|{E_{u,v}}\|_{2}$ (4)
Figure 3: The architecture of BriLLM.
Training BriLLM involves constructing a dedicated neural network for each sequence sample, which is illustrated in Figure 4. The network connects input nodes sequentially, with all potential paths integrated into a final softmax layer that identifies the correct path via a 0-1 ground-truth vector. Cross-entropy loss optimizes this process.
Figure 4: The training network of BriLLM for one training sample .
4 Experiments
We released BriLLM-Chinese and BriLLM-English, with experimental details below.
Figure 5: The training loss.
Datasets
BriLLM-Chinese and BriLLM-English were trained on Chinese and English Wikipedia, respectively (each with over 100M tokens). Long sentences were truncated to 32 tokens, with a 4,000-token vocabulary for both languages.
Implementation Details.
BriLLM is implemented in PyTorch, using sine-cosine positional encoding, GeLU activation, and cross-entropy loss for next-token prediction. Nodes have dimension $d_{node}$ = 32. Training used the AdamW optimizer ($\beta_{1}$= 0.9, $\beta_{2}$ = 0.999, $\epsilon=10^{-8}$) on 8 NVIDIA A800 GPUs for 1.5k steps. The theoretical model size is 512 + 4000 $\times 4000\times$ (32 $\times$ 32 + 32) $\approx$ 16B parameters, with training loss shown in Figure 5.
Sparse Training
BriLLM leverages sparse token co-occurrence: most bigrams are low-frequency or absent, allowing shared parameters for inactive edges. Low-frequency bigrams use a fixed, non-updated matrix, reducing model size to 2B (Chinese) and 1B (English)—13.0% and 5.7% of the original size, respectively (Table 2). This reduces parameters by  90% while accelerating training.
Table 2: Model sizes before and after sparse training.
BriLLM-Chinese BriLLM-English
original 16.90B 16.90B
sparse 2.19B 0.96B
ratio 13.0% 5.7%
Complexity
Forward computation complexity is O($L\cdot n\cdot d_{node}^{2}$), where $L$ is sequence length, $n$ is vocabulary size, and $d_{node}$ is node dimension.
Case Study
Tables  3 and  4 show decoding results for training and test samples in Chinese and English, respectively.
Input Completion
Training samples
《幽明录》,亦作 《幽明录》,亦作《幽冥录》、《我
《罗马》描述了 《罗马》描述了古罗马从共和国走下
众多哥萨克领导人开始 众多哥萨克领导人开始改变阵营运动
阿根廷探戈是起源于 阿根廷探戈是起源于阿根廷或乌拉圭
惠安第一中学是 惠安第一中学是位于福建省泉州市天
提琴本泛指 提琴本泛指有琴杆,需一手持杆另一
塞人定义为 塞人定义为阿塞拜疆人或南高加索地
袁乃宽旧居是 袁乃宽旧居是清朝及中华民国军事情
飞翔公园站是 飞翔公园站是广州地铁二号线的一。
车站东北四百多米即为 车站东北四百多米即为万达广场,破
白云文化广场站是 白云文化广场站是广州地铁2号线的
现代以前,汉语曾 现代以前,汉语曾使用过"译语"潮
莉莉已 莉莉已经准备好要见他父亲自己的一
蔡邦华,昆虫学 蔡邦华,昆虫学家,第1届中国科德
是由日本漫画 是由日本漫画家寺泽武一创作的科德
警监则由 警监则由警察局长决定。另外两岸气
,他的小提琴 ,他的小提琴演奏技巧备受关注。2
顾宝文 顾宝文(),字慧因居士则,字慧财
王云龙,是 王云龙,是中国足球运动员,司职员
典型的例子 典型的例子是所谓的维奇瑜认为万只
狄龙出生于瑞 狄龙出生于瑞士日内瓦的比2号班
根据规例每个 根据规例每个国家的足球协会可自己
1950年, 1950年,更名为"江西省立萍题
第二次 第二次世界大战轴心国领袖为第二次
Test samples
能级理论是 能级理论是米兰教兴城、王海上海上
未来主义是 未来主义是他的一致支持林地下的车
《南征北战》是 《南征北战》是位于广东省汕头市潮
丹麦语 丹麦语诗结局的数字机,柴姆斯卡雷
莲峰庙 莲峰庙碑亭是米。",设立为那亚州
他也不认为 他也不认为一个地区()是一个地区
卓越工程师 卓越工程师评量大陆的固的选择权—
群众只能够 群众只能够喷嘴能随即在宗,每年去
晚些时候 晚些时候阮惠安岭林斯·罗力发的第
他是 他是日返自行车特的一部,但没有的
Table 3: Case study of BriLLM-Chinese decoding results.
Input Completion
Training samples
In frogs, the hind legs are larger In frogs, the hind legs are larger than taxation arrangements and terms, misconstd Paris Academy members of Portals
The requirement for the Sun angle was The requirement for the Sun angle was arguments from Intr proposed: documentary directed by employing hundreds reduced by employe 11 September 1972
The English biologist Thomas Henry Huxley The English biologist Thomas Henry Huxley coined World C that ADE XaZul 30 Ars lead singular shipb more smaller im
Physicist Richard Feynman was noted for facility Physicist Richard Feynman was noted for facility in him increasingly holding six countries, misconstd atomic freedom before
Elements heavier than iron were Elements heavier than iron were retreatywriter 10th worked (ital magnitude, misconstd atomic Music freedom
Typically, when an algorithm is associated with Typically, when an algorithm is associated with Achill declaraus, misconceptions presented at Irraditional emotunday Prich
Plants are used as herbs Plants are used as herbs and Earth Day of Portals working on recent years of Portals working on recent genocots only marked serious risk that
The term vestibular The term vestibular at Texas variable Spec struggathological ideal remains the division of value of value cannot be supern2
Knight’s criticism greatly damaged van Knight’s criticism greatly damaged vanand soon to: examples are ’to looked identity said to: accounts reduced by employe
Atlas-Imperial, an American Atlas-Imperial, an American Advideo game), December with Achill declar between 2003, misconstd atomic freedom in
Test samples
The islands have The islands have been cultivated less than form of value and 1969 via the division of value, miscons lead to non-ane rock
The blue whale (Balaenoptera musculus) The blue whale (Balaenoptera musculus) order in him responsibility of Portals working on recent gene 11 September 197
The Vincent Price film, House of Wax The Vincent Price film, House of Waxi theorem approached the sequel strikend across the sequel strikend across
The Jewish Encyclopedia reports, In February The Jewish Encyclopedia reports, In February 11th worked in him increasingly holds reduced by employe 11 September 1972
The Bermuda Triangle The Bermuda Triangle, Azerbaijani official letters) markeditors), highest number of Portals working on recent years, misconcept of
Table 4: Case study of BriLLM-English decoding results.
5 Conclusion, Limitation and the Future
BriLLM introduces a transformative framework for language modeling, replacing attention-based architectures with brain-inspired dynamic signal propagation over a fully connected graph. By representing tokens as nodes and using energy dynamics to guide signal flow, BriLLM enables non-autoregressive generation, full node-level interpretability, and theoretically infinite $n$-gram modeling. Its design decouples model size from sequence length, optimizing resource use while mimicking neurocognitive processes like memory formation. This work challenges the dominance of attention mechanisms, offering a scalable, transparent alternative aligned with neural signaling principles.
Current limitations stem from constrained computational resources, with early checkpoints using moderate hyperparameters—yet performance already matches GPT-1 (Radford et al., 2018). A key distinction from traditional machine learning lies in SiFu’s support for concurrent multi-input/multi-output streams, whereas conventional models process a single input-output pair at a time. We envision an embodied BriLLM where multiple independent signal streams propagate simultaneously, guided by energy maximization—enabling parallel thought processes analogous to human multitasking, a capability beyond traditional single-stream architectures.
SiFu learning differs fundamentally from representation learning in deep learning. In BriLLM, the primary learnable parameter is the bias vector $b$, which lacks explicit representational intent; its role is to filter signal flows rather than encode semantic content. Even if interpreted as a weak form of node embedding, $b$ is far removed from the robust representational frameworks central to deep learning.
BriLLM’s current parameter count is ($n\times(1+d_{node}))^{2}+n^{2}\times d_{node}$ (with $n$ as token count and $d_{node}$ as node dimension), leading to quadratic growth. However, sparse training—using shared parameters for low-frequency token pairs—reduces size by  90% (Table 2), demonstrating potential for efficient scaling.
Training efficiency remains a challenge: BriLLM’s dynamic per-sample network construction hinders parallelization, and energy-based ranking over large vocabularies demands extensive computations. Approximate ranking—comparing candidate tokens only to relevant counterparts—could mitigate this, enabling parallel training of non-overlapping local networks.
Note that even though BriLLM theoretically supports infinite-gram language model without increasing model size, in practice, the model during training has to cover long enough input sequences, otherwise BriLLM decoding cannot give good enough sequence prediction beyond the training sample length. However, facilitating longer sequence prediction in terms of BriLLM simply depends on longer training without resizing the model itself.
BriLLM’s full interpretability enables innate multi-modal integration: nodes need not be limited to linguistic tokens but can represent units from diverse modalities. Critically, redefining nodes for new modalities does not require retraining from scratch—unlike LLMs, which rely on input/output alignment.
The initial signal vector (Eq. 1) is uniformly initialized, but we have validated stability with random initialization. Future work will explore signal shaping as a form of prompt engineering, potentially enabling more natural context specification than in-context learning in LLMs.
BriLLM’s performance at 1-2B parameters, comparable to 0.1B GPT-1, suggests a high threshold for LLM-like emergent capabilities. Additionally, supervised fine-tuning (SFT) methods for BriLLM remain underexplored, representing a key area for future research.
Table 5: Comparison of GPT-LLM and BriLLM.
LLM BriLLM
model size correlated to input context length independent
interpretability only in input & output all nodes throughout the model
multi-modal implementation limited to be joined from input/output all nodes throughout the model